### v2

# Cross-Cultural Expert-Level Art Critique Evaluation with Vision-Language Models

Abstract
Vision-Language Models (VLMs) excel at visual perception, yet their ability to interpret cultural meaning in art remains under-validated. However, cultural understanding and interpretability are often overlooked when evaluating these models. To overcome this limitation, this paper introduces a tri-tier evaluation framework for cross-cultural art-critique assessment. Tier I provides a series of automated metrics indicating cultural coverage. Tier II leverages theory-informed template-based scoring using a single primary judge across five evaluation dimensions (Coverage, Alignment, Depth, Accuracy, Quality), each rated on a 1–5 scale. Tier III then calibrates the aggregated scores from Tier II via isotonic regression. The proposed evaluation framework is validated with a large-scale experiment covering 15 different VLMs on 294 evaluation art-critique pairs spanning six different cultural traditions. Our findings reveal that (i) automated metrics are unreliable for cultural depth analysis, (ii) Western samples score higher than non-Western samples under our sampling and evaluation template, highlighting potential model biases, and (iii) VLMs exhibit a consistent performance gap, performing well in visual description but underperforming in cultural interpretation. Dataset and code are available at https://github.com/yha9806/VULCA-Framework.
Cross-Cultural Expert-Level Art Critique Evaluation with Vision-Language Models
Haorui Yu1, Ramon Ruiz-Dolz2, Xuehang Wen3, Fengrui Zhang4, and Qiufeng Yi5 1DJCAD, University of Dundee, United Kingdom 2Faculty of Science, Engineering and Business, University of Dundee, United Kingdom 3Hebei Academy of Fine Art, China 4Computer Science, Nanjing University, China 5School of Computer Science, University of Birmingham, United Kingdom
1 Introduction
Vision-Language Models (VLMs) achieve high scores on object-existence probing and VQA-style benchmarks (e.g., POPE Li et al. (2023); VQAv2 Goyal et al. (2017)). Yet these metrics mask a critical limitation: VLMs excel at identifying what is there but fail to interpret what it means. In this paper, we operationalize cultural understanding through five levels: L1 (visual perception), L2 (technical analysis), L3 (cultural symbolism), L4 (historical context), and L5 (philosophical aesthetics). Figure 1 illustrates this limitation, while a VLM correctly identifies “mountains and mist” (L1), it completely misses the Daoist philosophy of emptiness (L5) that gives the painting its cultural significance.
Input: Yuan dynasty landscape painting VLM: “Mountains in mist. Ink wash with varying tones.” Expert: “Sparse brushwork embodies yi (意); pine symbolizes scholarly integrity (L3); composition follows Ni Zan’s dry-brush tradition from Yuan literati school (L4); empty space reflects Daoist wu (无) philosophy (L5)…” Gap: VLM stops at L1–L2; Expert spans L3–L5
Figure 1: VLMs describe visual features (L1–L2) but miss cultural meaning (L3–L5). The expert critique explicitly covers symbolism (L3: pine = integrity), historical context (L4: Ni Zan, Yuan literati), and philosophy (L5: Daoist emptiness).
This limitation has real-world consequences, since VLMs are being widely integrated in many different use cases. Museums deploy VLM-powered guides, art education uses VLMs for critique assistance, and cultural heritage digitization relies on automated annotation. If VLMs mainly focus on superficial aspects of art, and systematically misinterpret non-Western art Cabello et al. (2023); Wu et al. (2025), these applications will perpetuate cultural misunderstanding at scale.
Existing benchmarks fail to capture cultural understanding for two major reasons. First, they assess perception without interpretation, benchmarks like MME Fu et al. (2025) and SEED-Bench-2 Li et al. (2024) test whether VLMs see objects (L1) but not whether they understand symbolism (L3) or philosophy (L5). Second, evaluation methods themselves can sometimes be unreliable. Dual-judge averaging suffers from systematic disagreement, we observe cross-judge Intraclass Correlation Coefficient (ICC) =0.50=-0.50, indicating that judges often operate on incompatible scales. Furthermore, automated metrics cannot substitute for human judgment, with correlations ranging only $r\in[0.27,0.53]$. In culturally-focused generative tasks, evaluation represents a major challenge. If the measuring instrument is unstable across judges and cultures, evaluation rankings become unstable and difficult to reproduce. The central challenge is therefore not only what to measure (e.g., cultural depth with L1-L5) but how to measure it with a controllable and validated instrument. Recent work Yu et al. (2025a, c) examined single cultures, finding significant VLM-expert divergence on Chinese painting critique. But do these failures reflect Chinese art complexity or systematic Western bias? Cross-cultural evaluation is needed to disentangle these factors.
To address these challenges, we introduce a Tri-Tier evaluation built on the L1–L5 cultural schema: Tier I involves automated coverage metrics, Tier II involves template-based single-judge scoring across five evaluation dimensions (i.e., Dimension Covering Rrate, Cultural Semantic Alignment, Critique Depth Score, Factual Accuracy Rate, Linguistic Quality Score), each rated on a 1-5 Likert scale, and Tier III addresses final calibration with human expert references. To avoid confusion, we refer to the evaluation stages with “Tier” and to the cultural understanding depth with “Level” (L1–L5).
This paper addresses three main research questions:
  1. RQ1:
    To what extent can VLMs assist in art critique and cultural understanding across L1-L5?
  2. RQ2:
    How large is the gap between Tier I and Tier II, and can calibration with human references close it?
  3. RQ3:
    Do VLMs exhibit systematic cultural biases in art critique tasks?
Our main contribution is therefore the evaluation framework itself, focusing on the cultural evaluation process rather than a new model architecture or training technique. The specific contribution of this paper is threefold: (1) a Tri-Tier framework combining automated metrics, single-judge scoring, and human reference calibration across 165 culture-specific dimensions (six evaluated traditions); (2) a set of evaluation assets including rubrics, prompts, and diagnostic scripts; and (3) an empirical analysis of 15 different VLMs revealing cultural understanding limitations and biases.
2 Related Work
2.1 VLM Evaluation Benchmarks
VLM evaluation has progressed through three generations. First-generation benchmarks (VQAv2 Goyal et al. (2017), GQA Hudson and Manning (2019), POPE Li et al. (2023), Winoground Thrush et al. (2022)) assess perceptual grounding at L1. Second-generation benchmarks (MME Fu et al. (2025), SEED-Bench-2 Li et al. (2024), MM-Vet Yu et al. (2024)) introduce multi-task evaluation, but “knowledge” remains encyclopedic rather than cultural.
Third-generation cultural probes reveal gaps: HallusionBench Guan et al. (2024), CulturalBench Chiu et al. (2025), and CulturalVQA Nayak et al. (2024) use multiple-choice QA rather than generative critique; GIMMICK Schneider et al. (2025) documents Western biases but employs recognition tasks rather than interpretation. None operationalize L1–L5 hierarchical understanding or provide frameworks for generative cultural critique.
2.2 Cultural AI and Bias Studies
Knowledge graph approaches such as WuMKG Wan et al. (2024) for Chinese culture and CulTi Yuan et al. (2026) for general cultural knowledge enable retrieval-based evaluation but do not cover generative evaluation of cultural elements. On the other hand, bias studies document disparities: Cabello et al. (2023) find Western-centric training correlates with non-Western performance degradation, and Wu et al. (2025) quantify demographic representation gaps. However, these studies measure classification accuracy rather than the depth of VLMs’ cultural understanding.
Prior work on VLM art critique reveals specific types of failure. Yu et al. (2025a) evaluated Chinese painting critique using 163 expert commentaries, finding VLM-expert divergence on qiyun shengdong (spirit resonance) interpretation. Yu et al. (2025c) documented “symbolic shortcuts”, i.e., VLMs correctly identify Western Christmas imagery but fail on non-Western festivals like Diwali. These single-culture studies cannot disentangle culture-specific challenges from systematic Western bias, highlighting the need for cross-cultural evaluation following a consistent methodology.
2.3 Computational Aesthetics
Style classification (e.g., WikiArt Saleh and Elgammal (2016) and OmniArt Strezoski and Worring (2018)), and affective computing-based approaches (e.g., ArtEmis Achlioptas et al. (2021)) mostly focus at L1-L2 levels without considering cultural interpretation in-depth. On the other hand, cultural modelling is still in its infancy, Zhang et al. (2024) apply Xie He’s Six Principles but only considering visual and technical aspects (L1-L2), and Jing (2023) theorizes yijing without conducting an evaluation of VLMs on this task. Our framework extends to cultural, historical, and philosophical aspects (i.e., L3-L5) through 165 culture-specific dimensions.
2.4 LLM-as-Judge Paradigm
Finally, when looking at LLM-as-judge approaches such as G-Eval Liu et al. (2023), MT-Bench Zheng et al. (2023), or Prometheus-Vision Lee et al. (2024), they provide scalable evaluation. However, we can observe that cultural critique tasks exhibits three systematic sources of error that dual-judge averaging cannot address: (1) strong judge disagreement (ICC =0.50=-0.50, Claude-Opus vs. GPT-5, $n=150$), (2) incompleteness (specifically, the absence of culture-specific dimensions for concepts such as Chinese wu or Japanese ma), and (3) distribution shift arising from Western-dominant training Tao et al. (2024).
Research Gap.
Table 1 compares the approach presented in this paper with prior work. We uniquely provide cross-cultural scope spanning 6 different cultural traditions with 165 culture-specific dimensions, hierarchical evaluation considering a deeper set of cultural aspects (i.e., L1-L5) rather than focusing on the surface (i.e.,L1-L2), a validated tri-tier framework with human-focused calibration that measurably improves judge-human alignment and evaluation stability, and systematic bias analysis with a quantified gap between Chinese and Western cultures.
Work Cultures Levels Validation
MME/SEED General L1 Auto only
CulturalBench 5 L3 (QA) Auto only
Zhang et al. (2024) Chinese L1–L2 None
Yu et al. (2025a) Chinese L1–L5 Expert
Ours 6 L1–L5 Tri-Tier
Table 1: Comparison with prior VLM cultural evaluation work.
3 Methodology
We propose a new evaluation methodology to assess cross-cultural art critique generation using a tri-tier framework (see Figure 2) that combines automated metrics, single-judge evaluation, and human-focused calibration.
Evaluation MethodologyInput: VLM Critique ccTier I: Auto Metrics(General Quality Indicators)DCRCSACDSLQSTier II: LLM Judge(Claude Opus 4.5)CovAlignDepthAccQualTier III: CalibrationIsotonic Reg. $g(\cdot)$$S_{\text{II}}^{*}=g(S_{\text{II}})$Cultural Understanding HierarchyL1 VisualL2 TechniqueL3 SymbolismL4 ContextL5 Philosophy  Primary path - - - Suplementary quality indicator
Figure 2: Vulca-Bench Evaluation Framework. Left: Evaluation Methodology where Tier I and Tier II (LLM judge) process the input in parallel. Tier II scores are calibrated considering human-expert references through Tier III producing $S_{\text{II}}^{*}$. Right: L1-L5 cultural understanding hierarchy.
3.1 Tri-Tier Framework
The evaluation framework processes a VLM-generated critique $c$ for culture $k$ through three tiers. Tier I and Tier II operate in parallel: Tier I computes four automated quality indicators, while Tier II produces template-based scores on five different cultural dimensions from a single primary judge. Then, Tier III calibrates the Tier II aggregate score to human expert ratings via isotonic regression.
Input/Output.
The framework receives as an input (i) a bilingual VLM-generated critique asked to cover L1-L5 levels of cultural understanding for a specific artwork image, and (ii) a culture tag. Then produces as an output (a) the calibrated aggregate score ($S_{\text{II}}^{*}$), (b) a template-based assessment covering five different dimensions, and (c) automated metrics-based risk indicators flagging potential types of failure (e.g., low cultural coverage, weak semantic alignment, or elevated template risk).
Tier I: Automated Metrics.
Tier I evaluation provides information about different aspects of an art critique based on simple heuristics. While they cannot replace a complete expert perspective, they provide a quantitative signal about potential limitations or weaknesses of a given critique. Therefore, we consider the following four automated metrics for a given VLM-generated critique $c$ and a specific culture $k$.
The Dimension Coverage Ratio (DCR, Eq. 1) measures the proportion of culture-specific dimensions $\mathcal{D}_{k}$ mentioned in $c$:
$\text{DCR}_{\text{auto}}(c,k)=\frac{|\mathcal{D}_{k}^{c}|}{|\mathcal{D}_{k}|}$ (1)
where $\mathcal{D}_{k}^{c}\subseteq\mathcal{D}_{k}$ are dimensions detected via keyword matching (e.g., $|\mathcal{D}_{\text{Chinese}}|=30$).
The Cultural Semantic Alignment (CSA, Eq. 2) measures TF-IDF cosine similarity with culture-specific vocabulary:
$\text{CSA}_{\text{auto}}(c,k)=\cos(\mathbf{v}_{c},\mathbf{v}_{k})$ (2)
where $v_{c}$ and $v_{k}$ are the TF-IDF vector representations of the critique and the culture-specific vocabulary respectively.
The Critique Depth Score (CDS, Eq. 3) measures whether a critique addresses deeper cultural understanding levels. We assign increasing weights to higher levels, as L3–L5 (symbolism, context, philosophy) require genuine cultural insight rather than surface description:
$\text{CDS}_{\text{auto}}(c)=\sum_{\ell=1}^{5}w_{\ell}\cdot\mathbb{1}\text{L}\ell\text{ covered}]$ (3)
where $\mathbb{1}[\cdot]$ is the indicator function (1 if level $\ell$ is detected via keyword matching, 0 otherwise), and $w_{\ell}=\ell/15$ assigns weights L1:1/15, L2:2/15, …, L5:5/15. A critique covering all five levels will achieve CDS=1.0 but one limited to L1–L2 will obtain a score of only 0.2, penalising critiques that focus on visual and technical aspects instead of providing deeper cultural understanding.
Finally, the Linguistic Quality Score (LQS, Eq. 4) detects fluent-but-shallow responses by combining length adequacy with sentence complexity:
$\text{LQS}_{\text{auto}}(c)=\min\left(1,\frac{|c|}{L_{\max}}\right)\cdot\frac{n_{\text{sent}}}{n_{\text{sent}}+\epsilon}$ (4)
where $|c|$ is the character count of critique $c$, $L_{\max}=2000$ is the expected length for an expert critique, $n_{\text{sent}}$ is the number of sentences, and $\epsilon=3$ is a smoothing constant. The first term rewards adequate length and the second term rewards multi-sentence structure. A long, single-sentence response will scores lower, flagging a potential template-style output.
Once calculated, all metrics are rescaled to $[1,5]$ so that they can be compared directly and aggregated within the same scale.
Tier II: Judge Scoring.
An independent judge (e.g., an LLM) scores each VLM critique following a template addressing five dimensions: Coverage (L1-L5 breadth), Alignment (culture-specificity), Depth (L3-L5 interpretation quality), Accuracy (factual correctness), and Quality (coherence). Each dimension receives a 1-5 rating which is then aggregated as their mean:
$S_{\text{II}}=\frac{1}{5}\sum_{i=1}^{5}d_{i}$ (5)
We adopt a single-judge design to ensure same-scale outputs. The LLM judge evaluates a critique based on the five-dimension template, then scores are calibrated via Tier III isotonic regression.
Tier III (Calibration).
Tier III calibration aligns the aggregate $S_{\text{II}}$ scores with human judgements via isotonic regression Niculescu-Mizil and Caruana (2005):
$S_{\text{II}}^{*}=\arg\min_{g\in\mathcal{G}}\sum_{i}(g(S_{\text{II}}^{(i)})-S_{\text{h}}^{(i)})^{2}$ (6)
where $\mathcal{G}$ is the set of monotonic functions and $S_{\text{h}}$ is the human score. Isotonic regression corrects systematic judge-human differences on the aggregate score. Dimension-specific scores, however, remain uncalibrated (Tier II) but are also taken into account for analytical purposes. Beyond reducing absolute error, calibration also enables more direct comparability across models and cultures by stabilising rankings under an explicit measurement scale based on human references.
4 Dataset
To validate our framework, we use the Vulca-Bench corpus Yu et al. (2026). The complete corpus contains 6,804 matched pairs across 7 cultural traditions. However, we exclude the Mural tradition (109 pairs) from evaluation due to an insufficient number of expert annotations for reliable Tier III calibration, resulting on a 6-culture evaluation subset of 6,695 image-critique pairs covering a total of 165 culture-specific dimensions (see Appendix A for a detailed list of the dimensions). From this subset, we extract 294 stratified samples with bilingual expert-written critiques which we use as evaluation samples (Eval.) for our experiments. On top of that, a disjoint 450-sample human-scored (HScored) subset of art critiques (298 training / 152 held-out test) serves as the gold-standard $S_{h}$ for fitting and validating the isotonic calibration function $g$ in Tier III. See Table 2 for a complete overview of our data distribution.
Each image-critique pair consists of (i) an artwork image, (ii) artist and title metadata, and (iii) an expert bilingual critique (Chinese: $\sim$500 characters; English: $\sim$300 words) annotated with culture-specific dimensions. Critiques explicitly identify L1–L5 layer coverage through dimension-specific tags (e.g., CN_L1_D1 for Chinese Visual Colour, see Appendix A), enabling the automatic computation of cultural coverage metrics.
Culture Dims Pairs Eval. HScored
Chinese 30 1,854 50 99
Western 25 4,270 50 97
Japanese 27 201 46 84
Korean 25 107 48 48
Islamic 28 165 50 52
Indian 30 98 50 70
Total 165 6,695 294 450
Table 2: Dataset distribution. Eval. column contains the sample distribution used for evaluating VLMs. HScored describes the distribution of human-scored art critiques.
Human-Scored Art Critiques.
The 450-sample of human-scored art critiques was annotated by 6 domain experts which assigned independent dual ratings to each art critique. Annotators scored based on our five Tier II dimensions (i.e., Coverage, Alignment, Depth, Accuracy, Quality) on a 1-5 scale. Inter-annotator agreement ($\kappa_{w}=0.39$) was fair Yu et al. (2025b), however, lower agreement on L3-L5 dimensions ($\kappa=0.35$) compared to L1-L2 ($\kappa=0.52$) evidenced the interpretive nature of this task, especially in deeper cultural dimensions such as philosophical aesthetics.
5 Experiments
5.1 Experimental Setup
We evaluate 15 VLMs on art critique generation considering our evaluation subset consisting of 294 samples. These models cover five providers including both closed and open weight models: GPT-5, GPT-5.2, GPT-5-mini, GPT-4o, GPT-4o-mini, Claude-Sonnet-4.5, Claude-Opus-4.5, Claude-Haiku-3.5, Gemini-2.5-Pro, Gemini-3-Pro, Qwen3-VL-235B, Qwen3-VL-Flash, Llama4-Scout, GLM-4V-Flash, and DeepSeek-VL2.
We compress images to $\leq$3.75MB, generate bilingual L1-L5 critiques using a unified prompt, and compute Tier I metrics. Claude Opus 4.5 (Tier II judge) then returns five 1-5 dimension scores (Coverage, Alignment, Depth, Accuracy, Quality) in JSON format. Scores are then calibrated using a 450-sample human subset (298 train / 152 held-out). Our default evaluation metric is the calibrated Tier III score ($S_{\text{primary}}=S_{\text{II}}^{*}$), although Tier II dimension-specific scores are also reported for analytical purposes.
5.2 Judge Selection
Model Provider Mean Std Lat. Tendency
GPT-4o OpenAI 4.52 0.17 2.9s Lenient
GPT-4o-mini OpenAI 4.31 0.16 1.6s Lenient
GPT-5 OpenAI 4.09 0.23 18.8s Moderate
Claude-Sonnet-4.5 Anthropic 3.99 0.16 14.5s Moderate
GPT-5.2 OpenAI 3.76 0.16 1.7s Strict
Claude-Haiku-3.5 Anthropic 3.76 0.37 7.6s High Var.
GPT-5-mini OpenAI 3.65 0.31 9.0s Strict
Claude-Opus-4.5 Anthropic 3.42 0.18 14.9s Strictest
Table 3: Single-judge comparison.
To select our Tier II judge, we compared 8 different models. As shown in Table 3, OpenAI models provided higher scores (mean 4.1-4.5) while Claude models were stricter (3.4-4.0). Furthermore, we also explored a dual-judge configuration, observing unreliable results where judges exhibited systematic scale mismatches. More specifically, we observed strong ICC disagreement in cross-judge scenarios ranging from -0.5 (Claude-Opus-4.5 + GPT-5, $n=150$) to 0.12 (Claude-Sonnet + GPT-5, $n=150$), all below our 0.6 threshold. In the end, single-judge design with isotonic calibration yielded stable scores aligned with human judgements. Therefore, we selected Claude Opus 4.5 for three main reasons: (1) stable rank discrimination, (2) consistent cultural gap direction, and (3) absence of self-favouritism (i.e., it does not favor its own outputs).
5.3 Calibration Validation
Dimension MAE MAE (cal)
Coverage 0.391 0.412
Alignment 0.511 0.482
Depth 0.537 0.486
Accuracy 0.534 0.552
Quality 0.461 0.431
Avg. 0.457 0.433
$\Delta$ -5.2%
Table 4: Human calibration results (n=152).
Table 4 shows the human calibration (Tier III) results. Isotonic calibration is done for each of the dimensions independently and then averaged as a final score. Although some dimensions (i.e., Coverage and Accuracy) do not improve with calibration, we observed an overall improvement of 5.1% in the final aggregated score. Our calibrated evaluation framework then produces this final aggregated score as its output ($S_{\text{II}}^{*}$), without focusing on specific dimensions.
6 Results and Analysis
6.1 Overall VLM Performance (RQ1)
Model $S_{\text{II}}^{*}$ Cov Align Depth Acc Qual TI
[…]
Gemini-3-Pro 3.78 4.50 4.44 4.51 4.48 4.49 3.37
Llama4-Scout 3.70 4.50 4.36 4.48 4.37 4.46 3.83
GPT-5 3.65 4.44 4.39 4.42 4.40 4.43 3.18
Claude-Opus-4.5 3.65 4.50 4.44 4.50 4.46 4.49 3.24
GPT-5.2 3.64 4.50 4.40 4.51 4.41 4.50 3.33
Qwen3-VL-Flash 3.53 4.50 4.54 4.48 4.56 4.48 3.51
GPT-5-mini 3.50 4.50 4.38 4.50 4.43 4.49 3.21
[…]
Std ($\sigma$) 0.03 0.13 0.10 0.13 0.08
Table 5: VLM $S_{\text{II}}^{*}$ performance with five Tier II dimensions (Cov=Coverage, Align=Alignment, Acc=Accuracy, Qual=Quality) and Tier I (TI) scores. The best results are in bold, the worst are underlined.
First, we analyse the overall VLM performance on our proposed evaluation framework. Table 5 shows a clear difference in performance when considering deeper cultural features among the 15 VLMs. Gemini-2.5-Pro (4.02) and Qwen3-VL-235B (3.99) are some of the best performing models, producing the most complete and deep critiques. On the other hand, Claude models remain competitive but do not achieve outstanding results (i.e., Sonnet 3.87, Opus 3.65). We can observe an overall spread of 0.78 points, revealing reasonable differences across different models. When looking at Tier II specific dimensions, we can see that Cultural Alignment and Factual Accuracy ($\sigma=0.13$) are the most discriminative dimension across models. For instance, we observe that DeepSeek-VL2 achieves reasonable Coverage but underperforms on Alignment/Accuracy, indicating weaker cultural grounding rather than surface-level fluency deficits.
We also observe that Tier II scores cluster in the 4.0-4.5 range ($\sigma=0.03$-$0.13$). This clustering reflects that (i) LLM judges have preference for polished language shaped by web-scale image-text pre-training, and (ii) ceiling effects in Coverage ($\sigma=0.03$), where most VLMs achieve >>80% dimension coverage and receive near-maximum scores. The discriminative power of our proposed framework lies in the higher-variance dimensions (i.e., Alignment and Accuracy, $\sigma=0.13$), which capture relevant L3-L5 cultural understanding features where VLMs show greater differences.
6.2 Tier I vs Tier II Gap (RQ2)
DCRCSACDSLQS1-10.5-0.500.50.5110.850.851.07-1.070.89-0.890.47-0.47Score Gap (TI - TII)
(a) Tier I-II gap. Positive = overestimate.
WE3.53.53.63.63.73.73.83.83.793.793.633.633.613.613.593.593.563.563.543.54Mean Score
(b) Cultural bias. Western highest.
Figure 3: (a) Tier I-II gap by metric. (b) Average performance scores by culture.
[…]
Table 6 compares Tier I metrics with the human-scored art critiques (n=450). All metrics show weak-to-moderate correlation ($r\in[0.27,0.53]$), confirming that they are not the best choice as individual evaluation metrics, but that can be useful as complementary signals. The information stemming from Tier I metrics can indicate critical limitations (e.g., superficial dimension coverage, lack of cultural depth) in an art critique, complementary to our calibrated judge output.
Metric Pearson $r$ Interpretation
[…]
Table 6: Tier I against human gold standard (n=450). All with $p<0.001$.
6.3 Cultural Bias (RQ3)
Figure 3 (b) shows that in general, Western art critiques were given higher scores by the LLM judge in our Tier II evaluation. More specifically, we observed a gap between scores attributed to Chinese and Western art of -0.17 (Cohen’s d=0.65d=-0.65, $p<0.001$).
When looking specifically at our models, we observe that eleven out of the fifteen favor Western art critiques, with GPT-4o showing strongest bias ($\Delta=-0.52$) while Claude-Opus-4.5 being the most neutral ($\Delta=0.00$). We interpret this as an observed trend under our sampling and evaluation template, rather than a definitive claim about intrinsic model biases.
7 Discussion
Table 7 illustrates the advantages of the proposed Tri-Tier framework for art critique evaluation. A high linguistic quality and richness can mask limited cultural understanding. The framework provides additional discriminant features that were missing on previous dual-judge setups by combining automated metrics (Tier I), five-dimensional template-based single-judge evaluation (Tier II), and evaluation calibration with human expert references (Tier III). Specifically, Tier I indicators flag fluent-but-shallow outputs where high Quality masks low cultural Alignment, as illustrated by the failure cases in Table 7.
Our findings have direct implications on the development of VLMs for cultural tasks. First, VLMs consistently perform well at visual-technical levels (L1–L2) but show significant variance at cultural-philosophical levels (L3–L5), with coverage drops up to 30%. This gap suggests that explicit hierarchical cultural information should be incorporated during pre-training, allowing models to understand deeper cultural aspects. Our error analysis shows that a majority of low-scored critiques mainly failed on Alignment and Accuracy dimensions, confirming a lack of cultural understanding as their main limitation. Furthermore, the low Tier I-II agreement (ICC <0.5<0.5) highlights that automated metrics cannot be a substitute. The case study in Table 7 further supports that specialized cultural modules may be necessary to complement VLMs in culture-specific symbolism understanding and interpretation.
Table 7: VLM critique excerpts illustrating success vs. failure patterns. Success cases demonstrate L3-L5 cultural depth with technique terminology and historical context, while failure cases remain at surface-level description despite fluent language.
Image Artwork Q Critique Excerpt Analysis
[Uncaptioned image] Wang Jian
Landscape Album
96 “…rocks rendered with pima cun (hemp-fiber strokes), combined with short pima and jiesuo cun…embodying the Song-Yuan landscape tradition Success
L2: technique terms
L4: historical lineage
[Uncaptioned image] Michelangelo
Creation of Adam
99 “…this ‘near-touch yet not quite touching’ moment becomes the dramatic climax, symbolizing life not yet fully awakened” Success
L3: symbolic reading
L5: theological depth
[Uncaptioned image] Shen Zhou
Farewell at Jingjiang
49 “This work by Ming dynasty Shen Zhou uses ink technique…composition is reasonable, brushwork is exquisite…overall artistic conception is profound” Failure
Template: generic praise
✗ No specific technique
[Uncaptioned image] E.A. Hornel
Springtime
67 “Color relationships function on multiple levels…careful balance of forms establishes visual equilibrium” Failure
L1 only: visual description
✗ No Glasgow School context
8 Conclusion
We introduced a Tri-Tier framework for cross-cultural art critique evaluation built on an L1-L5 cultural-depth schema. Tier I provides four lightweight, reference-free automated metrics for lower-bound assessment, Tier II applies template-based scoring with a single primary judge across five dimensions (Coverage, Alignment, Depth, Accuracy, Quality), and Tier III calibrates the aggregated Tier II scores to human expert ratings via isotonic regression.
We evaluated 15 VLMs across six different cultural traditions on a 294-sample evaluation set, and calibrated with a 450-sample subset containing human-scored art critiques, observing significant variations of performance in cultural understanding across five different dimensions. The dimensions with major variability (i.e., Alignment and Accuracy) highlight the limitations of VLMs as art critique generators, indicating that cultural grounding and factual accuracy still represent major challenges in culture-related tasks. Furthermore, we also observed a biased tendency between Western and Chinese art critiques, assigning higher scores to Western samples and being consistent across judge models. Finally, we emphasize that the proposed evaluation framework is designed to assist, not replace, expert interpretation in culturally sensitive settings.
Our study, therefore, provides relevant results pointing to some of the main limitations of current VLMs in cultural tasks. Future work will focus on implementing these improvements, seeking to deepen cultural understanding and reduce biases between different cultures.
Limitations
Cultural Coverage.
Our evaluation covers six cultural traditions (6,695 matched pairs: Western 4,270; Chinese 1,854; Japanese 201; Korean 107; Islamic 165; and Indian 98). Chinese and Western samples dominate the distribution, and this imbalance is reflected in our calibration experiments: the held-out test set contains fewer Korean ($n=16$) and Islamic ($n=18$) samples than Chinese or Western ($n=33$ each), and our analysis shows calibration degrades for Islamic ($6.5\%$) and Indian ($6.4\%$).
Language Coverage.
Our bilingual (Chinese–English) design reflects practical constraints. Expert annotators were only available for this language pair, and Chinese preserves key aesthetic terminology (qiyun, yijing) essential for keyword-based Tier I metrics. However, this choice may introduce translation loss for non-Chinese traditions: Japanese (wabi-sabi, mono no aware), Korean (jeong, heung), Islamic (Arabic calligraphic concepts), and Indian (rasa, bhava) art possess native terminologies that English romanization can only capture partially. We mitigated this by including romanized terms with diacritics where possible and training annotators on culture-specific vocabulary. However, extending the framework to native-language critiques (Japanese, Korean, Arabic, Hindi/Sanskrit) represents an important future direction.
Task Subjectivity.
The Tri-Tier framework depends on a single judge (Claude Opus 4.5) for Tier II scoring and on a 450-sample human-scored subset (298 train / 152 held-out) for isotonic calibration. Cross-cultural art critique evaluation, however, is inherently challenging because L3–L5 dimensions require interpreting philosophical concepts (qiyun, wabi-sabi, rasa) where legitimate expert disagreement reflects the interpretive nature of the task.
Few-Shot Prompting.
[…]
Appendix A 5-Level Framework Dimensions
Table 8 summarizes the 165 culture-specific dimensions across six traditions. Each culture has L1–L5 levels; dimensions marked with $\star$ are mandatory for quality validation ($\geq$70% coverage threshold).
Culture L1 L2 L3 L4 L5 Total
Chinese (CN) 6 6 6 6 6 30
Western (WE) 6 6 5 4 4 25
Japanese (JP) 6 6 5 5 5 27
Korean (KR) 5 5 5 5 5 25
Islamic (IS) 5 5 6 6 6 28
Indian (IN) 6 6 6 6 6 30
Total 34 34 33 32 32 165
Table 8: Culture-specific dimension counts by level.
Chinese (CN, 30 dims).
L1 (Visual): colour usage, composition, brushstroke texture, spatial levels, line expression, negative space (留白). L2 (Technical): cun texture strokes (皴法), brush techniques, ink gradation, colouring, brush dynamics, inscriptions/seals. L3 (Symbolic): imagery interpretation, symbolic meaning, cultural motifs, poetry-painting unity, literati aesthetics, aesthetic ideals. L4 (Historical): dynastic style, artist background, lineage, school characteristics, innovation, legacy. L5 (Philosophical): yijing (意境, artistic conception), qiyun shengdong (气韵生动, spirit resonance), heaven-human unity, aesthetic taste, form-spirit relation, void-solid interplay.
Western (WE, 25 dims).
L1: colour palette, line quality, composition, perspective, texture, chiaroscuro. L2: brushwork, colouring technique, medium/materials, subject/genre, style/school, technical mastery. L3: iconography, symbolism, signature/inscription, patronage/function, literary references. L4: period evolution, artist biography, influence/legacy, provenance. L5: aesthetic ideals, sublime/emotion, cross-cultural influence, artistic value.
Japanese (JP, 27 dims).
L1: colour, line, composition, spatial treatment, texture, light/shadow. L2: brushwork, colouring, subject matter, school/style, period style, innovation. L3: iconographic symbols, aesthetic concepts (wabi-sabi, mono no aware), inscriptions/seals, mounting format, aesthetic ideals. L4: period characteristics, school evolution, artistic lineage, social context, artistic exchange. L5: wabi-sabi (imperfect beauty), mono no aware (pathos of things), yūgen (profound grace), narrative aesthetics, artistic philosophy.
Korean (KR, 25 dims).
L1: colour usage, line expression, composition, spatial treatment, spirit resonance (기운). L2: brush techniques, ink methods, texture strokes, material properties, mounting format. L3: landscape imagery, Four Gentlemen, Confucian-Daoist spirit, minhwa folk painting, court painting. L4: dynastic periods (Goryeo/Joseon), school lineage, artists, Sino-Korean exchange, provenance. L5: literati aesthetics, jingyeong (true-view landscape), genre aesthetics, Korean identity, artistic value.
Islamic (IS, 28 dims).
L1: colour (lapis lazuli blue, vermilion), line quality, composition, spatial treatment, decorative patterns (arabesques). L2: negārgāri technique, pigment preparation, calligraphy (nasta‘līq), paper craft, preservation. L3: epic narrative (Shāhnāmeh), court life, religious themes, garden imagery (Chār Bāgh), architecture, natural symbols. L4: Persian schools (Herat, Tabriz, Shiraz), Mughal fusion, Ottoman tradition, Central Asian, dynastic patronage, manuscript tradition (kitābkhāneh). L5: calligraphic aesthetics, decorative philosophy, narrative aesthetics, Sufi spirituality, paradise vision, light symbolism.
Indian (IN, 30 dims).
L1: colour (mineral pigments, gold), line, composition, spatial, figure style, decorative elements. L2: miniature technique, pigment craft, paper (wasli), brushwork, preservation, signature. L3: Hindu iconography (Krishna, Radha), Rāgamālā (musical modes), romance narrative, court life, natural symbols, mythology. L4: Mughal court, Rajasthani schools, Pahari style, Deccan tradition, folk art, Company period. L5: Rasa theory (nine aesthetic emotions), Darshan tradition (sacred seeing), Bhakti devotion, decorative aesthetics, narrative aesthetics, spiritual realm.